# Determination of pairing matrix elements from average single particle level densities

### v1

@@ Abstract
A simple and efficient method to treat nuclear pairing correlations within a simple Hartree-Fock–plus-BCS description is proposed and discussed. It relies on the fact that the intensity of pairing correlations depends crucially on level densities around the Fermi surface ($\rho(e_{F})$) and that any fitting of nuclear energies as functions of the nucleon numbers is akin of a semi-classical average, smoothing out their quantal structure. A particular attention has been paid to two points generally ignored in previous similar approaches. One is a correction advocated by Möller and Nix [Nucl. Phys. A 536, 20 (1992)] taking into account the fact that the data included into the fit correspond to $\rho(e_{F})$ values systematically lower than average. The second is due to a systematic overestimation of the proton sp level density at the Fermi surface resulting from the local Slater approximation of the Coulomb terms in use in most microscopic descriptions. Our approach is validated by the agreement with data of corresponding calculated moments of inertia of well and rigidly deformed rare-earth nuclei, evaluated according to the Inglis-Belyaev ansatz with some crude Thouless-Valatin corrections. Indeed, the agreement which is found, is at least of the same quality as what results from a specific fit of the pairing intensities to these particular pieces of data. While this approach is currently limited to the very simple seniority force pairing treatment, it may serve as a starting point to define pairing residual interactions from averaged odd-even mass differences data, using merely average sp level densities associated to calculated canonical basis.
@@ II.1 Overview of the approach
Assuming that we know the smooth behaviour of the average neutron and proton pairing gaps with $N$ and $Z$, we devise here an approach to get the corresponding smooth evolution of pairing matrix elements $V_{q}$ (where $q$ stands for the charge states) averaged over a given sp valence space being contained in the $[\lambda_{q}-\Omega,\lambda_{q}+\Omega]$ range where $\lambda_{q}$ is the Fermi energy to be defined later, while the above energy interval (spanning a $2\Omega$ energy range) characterizes the domain of sp states (of the canonical basis) active in the BCS treatment. In this work we will take $\Omega=6$ MeV.
@@ II.3 Empirical vs effective-pairing gaps
The next crucial ingredient to the approach are the pairing gaps extracted from the data on $\delta E$, entering equation (1) or (3). At this point, it is important to differentiate between empirical pairing gaps as used e.g. in Refs. [9, 10] and those to be employed in any fitting approach such as ours. The latter are referred to as effective interaction pairing gaps in [11].
@@ II.4 Effective pairing gaps in the proton case
It is not as simple as to perform calculations with these $V_{p}^{0}$ values taking exactly into account the Coulomb exchange calculations to get $\Delta_{p}^{exact}$ and then determine $R_{p}$ as defined in eq (10) to correct for the Möller-Nix parameter. It turns out that the value of $R_{p}$ depends on the chosen value of $V_{p}^{0}$ or in effect of the degree of pairing correlations as a monotonically increasing function. This could be expected since the more pairing correlations are present, the closer to a smoothed-out level density distribution one would get, making it more and more closer to the one present in infinite nuclear matter corresponding to the Slater approximation.
Let us assume, as it will be discussed in Appendix A, that we have at our disposal a universal formula relating $R_{p}$ with a quantity representing the degree of pairing correlations. Here, we consider the so-called pair condensation energy $E_{cond}^{p}$ for protons (the absolute value of the part of the HF+BCS energy which involves the abnormal density), given by
$E_{cond}^{p}=\frac{\Delta_{p}^{2}}{V_{p}}.$ (12)
Having at hand the function $R_{p}(E^{p}_{cond})$, we will proceed as follows.
We perform HF+BCS calculations with Coulomb à la Slater with the matrix elements $V_{p}^{0}$. We then obtain BCS pairing gaps $\Delta_{q}$ and in particular the proton gap $\Delta_{p}(V_{p}^{0})$ which combined with $V_{p}^{0}$ will provide us with $E_{cond}^{p}(V_{p}^{0})$ and the associated value $R_{p}^{0}$ of the gap ratio defined in eq. (10). It is clear that this ratio depends on the retained value for the initial proton pairing matrix element $V_{p}^{0}$. The necessity for an iterative determination of the correct ratio $R_{p}$ consistent with a corresponding matrix element $V_{p}$ should, in principle, be advocated.
$r_{corr}=r\>R_{p}^{0}$ (13)
without reiterative procedure. We get the final value $V_{p}^{1}$ of the proton average pairing matrix element as done in the first stage of this calculation for a given nucleus an a given interaction, using now the corrected value $r_{corr}$ of the Möller-Nix parameter for the estimation process.
As detailed in Appendix A, such a determination of the values of $R_{p}$ has been achieved for a restricted sample (with respect to the 45 nuclei listed in Table 1) of 19 nuclei in the case of SIII (17 nuclei for SLy4 and SkM*) Specifically, we have excluded nuclei exhibiting large large sp energy gaps leading to artifically low BCS pairing gap, in view of the well-known deficiency of the BCS approximation in such weak pairing regimes (see e.g. Ref [23]).
Table 3: The average Fermi level densities for neutrons and protons (in MeV-1) are reported in columns 4 and 5, respectively, for the 176Yb and 240Pu nuclei in the ground-state deformation characterized by its quadrupole moment $Q_{20}$ (in barns). Columns 6 and 7 give the corresponding fitted fitted values (in MeV) of the average matrix elements $V_{q}$ (with $M=2$, $\beta=1.2$). These values are generated from HF+BCS calculations using different $G_{q}$ values (equal for neutrons and protons) as reported in column 2.
@@
III Technical details
III.1 Generation of sp levels
We have performed HF+BCS calculations for the deformed nuclear ground-states using three Skyrme parametrizations, namely the SIII [17], SkM* [18] and SLy4 [19], for the strong interaction part of the particle-hole interaction.
$V_{q}=\langle{i_{q}\bar{i_{q}}}|\widehat{v}_{res}\Big{(}|{j_{q}\bar{j_{q}}\rangle-|\bar{j}_{q}j_{q}}\rangle\Big{)}.$ (14)
$V_{q}=\frac{G_{q}}{11+N_{q}},$ (15)
and in what follows $G_{q}$ will be referred to as the pairing strength.
III.2 Choice of coefficients to determine average level density
To determine optimal values of $M$ and $\beta$, we performed fits of pairing matrix elements with $\Omega_{1}=\Omega_{2}=6$ MeV for two rare earth (166Dy and 176Yb) and two actinide (240U and 252No) nuclei. Figure 2 shows the variation of neutron (top panels) and proton (bottom panels) pairing matrix elements with $M$ and $\beta$ for these nuclei.
IV Results
where $N$ is the total number of sample nuclei. The r.m.s deviation have been analysed for 23 nuclei around rare-earth region with all three Skyrme parametrizations and 12 (13) actinide nuclei with SIII and SLy4 (SkM*) separately. The $\chi_{\mathcal{J}}$ are tabulated in the top part of Table 4.
Table 4: Weighted root-mean-square deviations $\chi_{\mathcal{J}}$ (in units of $\hbar^{2}/\mbox{MeV}$), $10^{4}\>\chi_{\mathcal{J}}^{A}$ (in units of $\hbar^{2}/\mbox{MeV}$) and $10^{2}\>\chi_{\mathcal{J}}^{exp}$ as defined in equations (16), (17) and (18 between calculated and experimentally defined MoI obtained with the three Skryme parametrisations.
SIII SkM* SLy4
$\chi_{\mathcal{J}}$ Rare earth 1.769 2.706 2.940
Actinide 4.582 3.394 3.498
$\chi_{\mathcal{J}}^{A}$ Rare earth 3.515 5.080 5.883
Actinide 4.844 3.588 3.621
$\chi_{\mathcal{J}}^{exp}$ Rare earth 4.548 7.230 7.551
Actinide 6.671 5.168 5.248
Table 5: Uncertainty ranges for the moment of inertia of some nuclei $\Delta_{J}^{A}$ and $\Delta_{J}^{exp}$ as defined in the text.
62 98 40.765 42.373 1.658 1.854
68 100 36.297 37.592 1.798 1.651
72 106 31.988 32.196 1.980 1.455
92 140 63.879 63.061 4.243 4.261
94 150 63.423 67.873 4.615 4.231
102 150 67.468 64.655 4.870 4.501
V Conclusion
The test of our method has consisted in using the so-determined average pairing matrix elements (with three different Skyrme force pararetrisations) to compare within an Inglis Belyaev approach (plus some approximate Thouless-Valatin correction) MoI of about forty well and rigidly deformed rare-earth and actinide nuclei, with what is deduced from the experimental energies of their first $2^{+}$ levels. The good quality of the theoretical estimate of these experimental MoI has been assessed to be about the same as what had been obtained within the same Hartree-Fock plus seniority BCS approach, in a particular case (for rare earth nuclei and using the Skyrme SIII parametrisation) by a direct fit of these MoI.
Acknowledgements.
Appendix A Reduction factor for the Moller-Nix parameter in the case of protons
We performed both exact Coulomb and Slater approximation calculations for a series of nuclei using the pairing matrix elements listed in Table 6. The obtained ratio $R_{p}$ of the BCS proton gaps from all three Skyrme parametrisations considered herein, are then plotted in Figure 4 as a function of proton condensation energy $E_{cond}^{p}$.
@@ Appendix A Reduction factor for the Moller-Nix parameter in the case of protons
$R_{p}=0.0181\>E_{cond}^{p}+0.781.$ (19)
This equation allows for an estimation of the reduction factor to the Möller-Nix parameter for any Skyrme parametrisation and at a given initial pairing matrix element $V_{p}^{0}$. Multiplying the initial Möller-Nix parameter $r=4.8$ with the reduction factor, one then obtains the proton pairing gap to be utilized in the estimation of pairing matrix element via equation (1).
Let us discuss finally the influence of the choice of the initial pairing matrix element $V_{p}^{0}$ on the retained value for the gap ratio and consequently on the renormalization of the Möller-Nix proton gap values to be used in the fit (see SubSection II.4). Upon calculating the ground state deformation solutions of our selected nuclei within the HF+BCS approach (using the seniority force model) we get proton pairing condensation energies. They are, as already noted, dependent upon the choice of the average pairing matrix elements in use in these calculations. From them, using equation 13 we should define new values of the proton gaps from which using the uniform gap method and the sp spectra, we could generate new proton pairing matrix elements $V^{1}_{p}$. And we could iterate this process to get a convergence of the pairing matrix elements.
The question now is how much change from $V_{p}^{0}$ could we expect for these new matrix elements $V^{1}_{p}$. To estimate that, we assume that we have chosen reasonable values of the original $V_{p}^{0}$, i.e. quantified, for instance, by a deviation of the resulting $E_{cond}$ from the value obtained through a converged solution of this iterative process by no more than ± $\pm\>20\%$. We see from Figure 4 that such an interval for $E_{cond}$ corresponds to an interval for $R_{p}$ (and thus on the proton gaps used in the fit) of ± $\sim\pm\>2\%$. From the approximation (shown to be rather good, see Figure 1) of equation 3 concerning the relation between $V_{p}$ and $\Delta_{p}$, we get readily
with $x=\frac{\Omega}{\Delta_{p}}$. Using the following values $\Omega=6$ MeV and $\Delta_{p}=1$ MeV, one obtains an uncertainty on $V_{p}$ within the ± $\pm\>0.8\%$ range.
A specific convergence study has been performed for the two deformed nuclei considered in Table 3 (176Yb and 240Pu) for the three values of the proton pairing intensity parameter $G_{p}=16,19,22$ MeV. One sees on Table 7 that the value of the proton matrix element $V^{0}_{p}$ is converged at the keV level at the second or third iteration. Similarly, the mass quadrupole moment $Q_{20}$ is converged at the fm2 level already at the third iteration, even sometimes at the second. The other lesson comparing the starting value of $V^{0}_{p}$ and the converged one, is that one may hint that a reasonable range for $V^{0}_{p}$ values would lie in between $G_{p}=16$ and $G_{p}=19$ MeV. This constitutes a very simple preliminary study, for any given specific particle-hole interaction, allowing to define a priori for a global study some value of $V^{0}_{p}$ close to a level of about $1\%$ to what an iterative process would produce.
Figure 4: Ratio of the proton BCS gap between exact Coulomb and Slater approximation calculations $R_{p}$ as a function of the proton pairing condensation energy $E_{cond}$. All energies are given in MeV.

### v2

@@ Abstract
A simple and efficient method to treat nuclear pairing correlations within a simple Hartree-Fock–plus-BCS description is proposed and discussed. It relies on the fact that the intensity of pairing correlations depends crucially on level densities around the Fermi surface ($\rho(e_{F})$) and that any fitting of nuclear energies as functions of the nucleon numbers is akin of a semi-classical average, smoothing out their quantal structure. A particular attention has been paid to two points generally ignored in previous similar approaches. One is a correction advocated by Möller and Nix [Nucl. Phys. A 536, 20 (1992)] taking into account the fact that the data included into the fit correspond to $\rho(e_{F})$ values systematically lower than average. The second is due to a systematic overestimation of the proton sp level density at the Fermi surface resulting from the local Slater approximation of the Coulomb exchange terms in use in most microscopic descriptions. Our approach is validated by the agreement with data of corresponding calculated moments of inertia of well and rigidly deformed rare-earth nuclei, evaluated according to the Inglis-Belyaev ansatz with some crude Thouless-Valatin corrections. Indeed, the agreement which is found, is at least of the same quality as what results from a specific fit of the pairing intensities to these particular pieces of data. While this approach is currently limited to the very simple seniority force pairing treatment, it may serve as a starting point to define pairing residual interactions from averaged odd-even mass differences data, using merely average sp level densities associated to calculated canonical basis.
@@ II.1 Overview of the approach
Assuming that we know the smooth behaviour of the average neutron and proton pairing gaps with $N$ and $Z$, we devise here an approach to get the corresponding smooth evolution of pairing matrix elements $V_{q}$ (where $q$ stands for the charge states) averaged over a given sp valence space being contained in the $[\lambda_{q}-\Omega,\lambda_{q}+\Omega]$ range where $\lambda_{q}$ is the Fermi energy to be defined later, while the above energy interval (spanning a $2\Omega$ energy range) characterizes the domain of sp states (of the canonical basis) active in the BCS treatment.
@@ II.3 Empirical vs effective-pairing gaps
The next crucial ingredients to the approach are the pairing gaps extracted from the data on $\delta E$, entering equation (1) or (3). At this point, it is important to differentiate between empirical pairing gaps as used e.g. in Refs. [9, 10] and those to be employed in any fitting approach such as ours. The latter are referred to as effective interaction pairing gaps in [11].
@@ II.4 Effective pairing gaps in the proton case
Table 3: The average Fermi level densities for neutrons and protons (in MeV-1) are reported in columns 4 and 5, respectively, for the 176Yb and 240Pu nuclei in the ground-state deformation characterized by its quadrupole moment $Q_{20}$ (in barns). Columns 6 and 7 give the corresponding fitted values (in MeV) of the average matrix elements $V_{q}$ (with $M=2$, $\beta=1.2$). These values are generated from HF+BCS calculations using different $G_{q}$ values (equal for neutrons and protons) as reported in column 2.
@@
It is not as simple as to perform calculations with these $V_{p}^{0}$ values taking exactly into account the Coulomb exchange calculations to get $\Delta_{p}^{exact}$ and then determine $R_{p}$ as defined in eq (10) to correct for the Möller-Nix parameter. It turns out that the value of $R_{p}$ depends on the chosen value of $V_{p}^{0}$ or in effect on the degree of pairing correlations as a monotonically increasing function. This could be expected since the more pairing correlations are present, the closer to a smoothed-out level density distribution one would get, making it more and more closer to the one present in infinite nuclear matter corresponding to the Slater approximation.
To quantify the degree of pairing correlations, we consider the so-called pair condensation energy $E_{cond}^{p}$ for protons (the absolute value of the part of the HF+BCS energy which involves the abnormal density), given by
$E_{cond}^{p}=\frac{\Delta_{p}^{2}}{V_{p}}.$ (12)
To determine the variation of $R_{p}$ as a function of $E_{cond}^{p}$, we consider a sample of 19 nuclei in the case of SIII (17 nuclei for SLy4 and SkM*) with respect to the 45 nuclei listed in Table 1. Specifically, we have excluded nuclei exhibiting large sp energy gaps leading to artifically low BCS pairing gap, in view of the well-known deficiency of the BCS approximation in such weak pairing regimes (see e.g. Ref [23]).
For these nuclei, we performed both exact Coulomb and Slater approximation calculations using the pairing matrix elements listed in Table 4. The obtained ratio $R_{p}$ of the BCS proton gaps from all three Skyrme parametrisations considered herein, are then plotted in Figure 2 as a function of the proton condensation energy $E_{cond}^{p}$.
@@
$R_{p}=0.0181\>E_{cond}^{p}+0.781$ (13)
where $E_{cond}^{p}$ is given in MeV. This equation allows for an estimation of the reduction factor to the Möller-Nix parameter for any Skyrme parametrisation and at a given initial pairing matrix element $V_{p}^{0}$. Multiplying the initial Möller-Nix parameter $r=4.8$ MeV with the reduction factor, one then obtains the proton pairing gap to be utilized in the estimation of pairing matrix element via equation (1).
@@
Having at hand the function $R_{p}(E^{p}_{cond})$, we will proceed as follows. We perform HF+BCS calculations with Coulomb à la Slater with the matrix elements $V_{p}^{0}$. We then obtain BCS pairing gaps $\Delta_{q}$ and in particular the proton gap $\Delta_{p}(V_{p}^{0})$ which combined with $V_{p}^{0}$ will provide us with $E_{cond}^{p}(V_{p}^{0})$ and the associated value $R_{p}^{0}$ of the gap ratio defined in eq. (10). It is clear that this ratio depends on the retained value for the initial proton pairing matrix element $V_{p}^{0}$. The necessity for an iterative determination of the correct ratio $R_{p}$ consistent with a corresponding matrix element $V_{p}$ should, in principle, be advocated.
$r_{corr}=r\>R_{p}^{0}$ (14)
without reiterative procedure. We get the final value $V_{p}^{1}$ of the proton average pairing matrix element as done in the first stage of this calculation for a given nucleus and a given interaction, using now the corrected value $r_{corr}$ of the Möller-Nix parameter for the estimation process.
III Technical details
III.1 Generation of sp levels
We have performed HF+BCS calculations for the deformed nuclear ground-states using three Skyrme parametrisations, namely the SIII [17], SkM* [18] and SLy4 [19], for the strong interaction part of the particle-hole interaction.
$V_{q}^{ij}=\langle{i_{q}\bar{i_{q}}}|\widehat{v}_{res}\Big{(}|{j_{q}\bar{j_{q}}\rangle-|\bar{j}_{q}j_{q}}\rangle\Big{)}$ (15)
for a given pair $(i,j)$ of the canonical basis states.
$V_{q}=\frac{G_{q}}{11+N_{q}},$ (16)
and in what follows $G_{q}$ will be referred to as the pairing strength.
Figure 2: Ratio of the proton BCS gap between exact Coulomb and Slater approximation calculations $R_{p}$ as a function of the proton pairing condensation energy $E_{cond}$ (given in MeV).
III.2 Choice of coefficients to determine average level density
To determine optimal values of $M$ and $\beta$, we performed fits of pairing matrix elements with $\Omega=6$ MeV for two rare earth (166Dy and 176Yb) and two actinide (240U and 252No) nuclei. Figure 3 shows the variation of neutron (top panels) and proton (bottom panels) pairing matrix elements with $M$ and $\beta$ for these nuclei.
IV Results
where $N$ is the total number of sample nuclei. The r.m.s deviation have been analysed for 23 nuclei around rare-earth region with all three Skyrme parametrisations and 12 (13) actinide nuclei with SIII and SLy4 (SkM*) separately. The $\chi_{\mathcal{J}}$ are tabulated in the top part of Table 5.
Table 5: Weighted root-mean-square deviations $\chi_{\mathcal{J}}$ (in units of $\hbar^{2}/\mbox{MeV}$), $10^{4}\>\chi_{\mathcal{J}}^{A}$ (in units of $\hbar^{2}/\mbox{MeV}$) and $10^{2}\>\chi_{\mathcal{J}}^{exp}$ as defined in equations (17), (18) and (19 between calculated and experimentally defined MoI obtained with the three Skryme parametrisations.
SIII SkM* SLy4
$\chi_{\mathcal{J}}$ Rare earth 1.769 2.706 2.940
Actinide 4.582 3.394 3.498
$\chi_{\mathcal{J}}^{A}$ Rare earth 3.515 5.080 5.883
Actinide 4.844 3.588 3.621
$\chi_{\mathcal{J}}^{exp}$ Rare earth 4.548 7.230 7.551
Actinide 6.671 5.168 5.248
Table 6: Uncertainty ranges for the moment of inertia (in units of 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV) of some nuclei $\Delta_{J}^{A}$ and $\Delta_{J}^{exp}$ as defined in the text.
62 98 40.765 42.373 1.658 1.854
68 100 36.297 37.592 1.798 1.651
72 106 31.988 32.196 1.980 1.455
92 140 63.879 63.061 4.243 4.261
94 150 63.423 67.873 4.615 4.231
102 150 67.468 64.655 4.870 4.501
V Conclusion
The test of our method has consisted in using the so-determined average pairing matrix elements (with three different Skyrme force parametrisations) to compare within an Inglis Belyaev approach (plus some approximate Thouless-Valatin correction) MoI of about forty well and rigidly deformed rare-earth and actinide nuclei, with what is deduced from the experimental energies of their first $2^{+}$ levels.
These sets of Skyrme parametrisations were chosen to allow us to assess the versatility of our pairing matrix elements estimation approach using different mean-field solutions. It was not our intention here to evaluate the quantal deficiency of the sp spectrum produced using any of the Skyrme parametrisations. In fact, we showed that the pairing matrix elements estimated using the proposed approach herein is not affected by the wrong reproduction of the sp energy gap. Therefore, this deficiency is irrelevant for a test of our approach. Consequently, we had to remove these nuclei in our samples because implementing them in our test would unduly affect the quality assessment of the proposed method.
From the remaining list of good sample nuclei, we found a rather good reproduction of the experimental MoI. This is especially so for the case of the Skyrme SIII parametrisation in the rare earth region where our present estimates give excellent agreement with results obtained from a direct fit to these MoI.
Acknowledgements.
Appendix A Treatment of the arbitrariness of the initial proton matrix element $V_{p}^{0}$
Here, we discuss how to circumvent the arbitrariness of the chosen initial pairing matrix element $V_{p}^{0}$. Upon calculating the ground state deformation solutions of our selected nuclei within the HF+BCS approach (using the seniority force model) we get proton pairing condensation energies. They are, as already noted, dependent upon the choice of the average pairing matrix elements in use in these calculations. From them, using equation (14) we should define new values of the proton gaps from which using the uniform gap method and the sp spectra, we could generate new proton pairing matrix elements $V^{1}_{p}$. And we could iterate this process to get a convergence of the pairing matrix elements.
The question now is how much change from $V_{p}^{0}$ could we expect for these new matrix elements $V^{1}_{p}$. To estimate that, we assume that we have chosen reasonable values of the original $V_{p}^{0}$, i.e. quantified, for instance, by a deviation of the resulting proton $E_{cond}$ from the value obtained through a converged solution of this iterative process by no more than ± $\pm\>20\%$. We see from Figure 2 that such an interval for $E_{cond}$ corresponds to an interval for $R_{p}$ (and thus on the proton gaps used in the fit) of ± $\sim\pm\>2\%$. From the approximation (shown to be rather good, see Figure 1) of equation 3 concerning the relation between $V_{p}$ and $\Delta_{p}$, we get readily
with $x=\Omega/\Delta_{p}$. Using the following values $\Omega=6$ MeV and $\Delta_{p}=1$ MeV, one obtains an uncertainty on $V_{p}$ within the ± $\pm\>0.8\%$ range.
However, a specific convergence study has been performed for the two deformed nuclei considered in Table 3 (176Yb and 240Pu) for the three values of the proton pairing intensity parameter $G_{p}=16,19,22$ MeV. One sees on Table 7 that the value of the proton matrix element $V_{p}$ is converged at the keV level at the second or third iteration. Similarly, the mass quadrupole moment $Q_{20}$ is converged at the fm2 level already at the third iteration, even sometimes at the second. The other lesson comparing the starting value of $V^{0}_{p}$ and the converged one, is that one may hint that a reasonable range for $V^{0}_{p}$ values would lie in between $G_{p}=16$ and $G_{p}=19$ MeV. This constitutes a very simple preliminary study, for any given specific particle-hole interaction, allowing to define a priori for a global study some value of $V^{0}_{p}$ close at a level of about $1\%$ to what an iterative process would produce.